# Distributed Split Computing Using Diffusive Metrics for UAV Swarms

### v1

@@ I Introduction
The escalating prominence of Unmanned Aerial Vehicle (UAV) swarms has opened up new horizons in applications spanning surveillance, disaster response, smart agriculture, and beyond. By coordinating large numbers of autonomous aerial units, system-level benefits such as enhanced coverage, redundancy, and adaptability can be realized. However, these advantages come hand-in-hand with steep technical challenges, particularly when it comes to designing and deploying machine learning (ML) services in vast swarms. Traditional methods that rely on a single, centralized server or controller become increasingly untenable as the number of nodes and geographic scale increase[1]. A centralized ground station, faces prohibitive communication overheads, latency bottlenecks, and reliability issues. Moreover, real-world UAV swarms are dynamic, meaning that nodes may join, leave, or rearrange themselves in the network. Such dynamic conditions only make the limitations of traditional centralized paradigms worse[2].
@@
Recent efforts to alleviate centralized burdens typically advocate a form of distributed operation by distributing inference or training tasks across multiple nodes[3]. However, they still often rely on global awareness or explicit synchronization. At smaller scales or in constrained network topologies, these partial-distribution schemes might suffice. Yet, as UAV swarms expand to hundreds or even thousands of nodes, it becomes infeasible to gather and coordinate global network information[4]. Additionally, heterogeneous node capabilities such as differences in on-board processing, sensing, and power resources demand an approach that can dynamically adjust to varying node states in near real-time. A node endowed with high computation power a time may deplete its battery and become less capable a moment later, or an agile UAV that was ideally placed to handle tasks at one moment could relocate out of range in the next. These variations highlight the necessity for a self-organizing approach that pushes intelligence and decision-making to the edges.
@@
Split learning and computing is particularly well-suited to address these challenges in large-scale UAV swarms[3]. Rather than transmitting raw data or entire models to a single aggregator, split learning and computing partitions the model into layers distributed among various nodes, thereby enabling partial inference at local levels. By transferring only the necessary, intermediate representations, nodes can reduce communication overhead while maintaining data privacy. However, effectively implementing split computing in UAV networks requires strategic decision-making about when and where to forward intermediate layer outputs[1, 3]. This is where centralized schemes fail to perform well for large networks. They attempt to collect global knowledge of the network’s topology, node resource levels, and data distributions which is impractical in large, dynamic swarms. As a result, model splits and partial inferences often end up bottlenecked by nodes that unknowingly become overloaded or unreachable.
@@
To further enhance the system’s adaptability under fluctuating workloads, we incorporate an early-exit mechanism that dynamically reduces the depth of inference based on real-time load conditions. This strategy allows nodes to truncate the model execution at predefined intermediate layers, producing acceptable outputs while avoiding the full computational cost of processing the entire model. By adjusting the exit point according to the current resource availability and task queue length, the system minimizes latency and prevents resource saturation. This approach aligns with the principles of early exiting [5] where intermediate outputs offer a controllable trade-off between computational effort and inference accuracy, making it particularly suitable for dynamic and resource-constrained UAV environments.
@@
The remainder of this paper is organized as follows. Section II reviews related work on distributed inference and task offloading in UAV networks. Section III formalizes the system model and problem definition, introducing the proposed diffusive metric and task allocation strategy. Section V presents the simulation setup and provides a detailed performance evaluation under various operating conditions. Finally, Section VI concludes the paper and outlines directions for future work.
@@ II Related Work
Recent research has explored distributed DNN inference and task offloading to address the challenges posed by limited resources and dynamic network conditions. In [6], MoDNN partitions DNN computations among mobile devices to balance workload and reduce communication overhead; its model partitioning strategy minimizes transmission delays while exploiting parallel processing, making it highly applicable to resource-constrained environments. Similarly, DISNET [7] introduces a micro-split framework that partitions deep neural network models both vertically and horizontally, enabling cooperative inference among heterogeneous IoT devices while reducing latency and energy consumption through adaptive workload allocation.
@@
On the split computing front, [8] propose a multi tier framework that employs deep reinforcement learning to optimize model splitting and resource allocation in UE–Edge–Cloud systems, decomposing complex optimization problems into sequential subproblems to balance computational and communication loads. In a similar vein, CoEdge [9] demonstrates that dynamic workload partitioning—formulated as a constrained optimization problem—can yield substantial latency speedups, with a fast approximation algorithm that adapts in real time to varying device capabilities and network conditions. Similarly for Split Learning, [10] develop an efficient parallel split learning (EPSL) framework for resource-constrained wireless edge networks. Their approach jointly optimizes subchannel allocation, transmit power, and model split point selection to minimize latency and communication overhead. [11] propose Edge-MSL, a split learning framework for mobile edge networks that leverages a contextual multi-armed bandit formulation to dynamically select the optimal split point and assign edge servers. By adapting the split location and offloading decisions based on real-time network conditions, their approach significantly reduces training latency and enhances convergence. On the other hand, [12] combines split learning with federated learning in UAV networks to address the trade-off between communication overhead and computational constraints, resulting in an architecture that achieves higher inference accuracy than standalone federated learning while reducing transmission costs compared to conventional split learning. Moreover, [13] propose a parallel split learning framework over wireless networks that integrates dynamic resource management with parallel task offloading. Their method exploits concurrent processing across heterogeneous edge devices to substantially reduce training latency while maintaining balanced workload distribution. [14] investigate a UAV-assisted split computing system where a UAV acts as a mobile edge server to support collaborative inference among IoT devices. By partitioning a DNN between on-device processing and UAV-based computation, their design reduces transmission delays and energy consumption.
@@
Task offloading approaches can also be useful for split learning approaches. For instance, [15] develops a UAV-based environmental monitoring system that employs on-board processing to minimize transmission delays and maintain low data freshness latency, thereby enhancing overall system capacity in remote areas. [16] proposes a cloud–MEC collaborative offloading scheme that integrates differentiated offloading decisions with service orchestration via an ODaS mechanism, significantly reducing both delay and energy consumption by balancing computational and communication loads. [17] addresses the challenge of offloading dependent tasks modeled as directed acyclic graphs using a branch soft actor-critic algorithm, which optimizes offloading, computation, and cooperation decisions under dynamic conditions to reduce latency and improve energy efficiency. More recently, [18] presents a multi-agent collaborative framework that jointly optimizes UAV trajectory planning and DAG task offloading via deep reinforcement learning, with their TD3-TT algorithm demonstrating further latency reductions and enhanced real-time performance. The FOCUS framework [19] leverages fog computing in UAV networks by jointly optimizing routing and computation. Its central SDN-based architecture and analytical approach achieve significant latency improvements by effectively integrating on-board processing with data forwarding. Lastly, [20] investigate task offloading in a UAV-assisted MEC network by leveraging a federated learning framework. In their work, a UAV equipped with MEC capabilities serves as a mobile server to assist resource-constrained UE in offloading computation-intensive tasks.
@@
III Problem Formulation and System Model
@@
III-A Task Computation Model
@@ III-B Communications
III-C Energy Consumption Model
@@
IV Distributed Task Allocation and Early-Exit Mechanism
@@
Nodes with powerful neighbors or central positions are more advantageous for task allocation. To identify such nodes, we aim to utilize graph centrality measures. However, these measures typically require a global view of the network, which conflicts with our distributed approach. Metrics like eigenvector centrality, which consider only a node’s neighbors, provide a feasible alternative. Eigenvector centrality for a node is calculated as:
$x_{v_{i}}={\frac{1}{\lambda}}\sum_{v_{j}\in M(v_{i})}x_{v_{j}}$ (8)
Here $M(v_{i})$ is the set of neighbors of $v_{i}$ and $\lambda$ is the largest eigenvalue of the network graph’s adjacency matrix. Eigenvector centrality is iteratively computed using power iteration: initial centralities are guessed and updated until convergence. Normalization is required at each step to prevent divergence, though this typically requires a global view. This iterative approach could work for distributed networks if divergence can be controlled.
To address these challenges, we propose using aggregated computation capability instead of eigenvector centrality. This metric measures how quickly a node can complete a job when it distributes the load among itself and its neighbors.
@@
We iteratively compute aggregated computation capability. The idea is if a load given to a node, if it were to distribute this load to its neighbors evenly, when this load would be completed. The aggregated computation capability $\phi_{i,t+1}$ of node $v_{i}$ at time $t+1$ is:
@@
where $M_{i}(t)$ is the neighbor set of node $i$, $F_{i}$ is computation capability of node $i$, $d^{tx}_{i,k}$ is the transmission delay between nodes $i$ and $k$, and $\phi_{i,t}$ is the aggregated computation capability of node $i$ and time $t$. The aggregated computation capability assumes that the node divides the work evenly ($|M(i)|+1$), computes its part, then it broadcasts the load to its neighbors. Since the neighbors can work concurrently, the task will be completed lastly by the slowest node with aggregated computation capability. Since this update requires only the neighbors of the nodes, it dynamically adapts the network conditions without need for global view.
IV-A Task Transfer Decision
Our framework requires each UAV to make autonomous decisions about whether to process a task locally or transfer it to a neighboring node. These decisions are based on comparative utilization levels between nodes, mediated by a threshold parameter $\gamma$ that controls the sensitivity of task offloading.
The utilization of node $i$ at time $t$ is calculated as the ratio of its current computational load to its aggregated computation capability:
@@ IV-A Task Transfer Decision
IV-B Early-Exit Mechanism
@@
This adaptive mechanism ensures that nodes under high load degrade computation gracefully, trading off accuracy to maintain responsiveness and system stability. It is important to note that the neural network architecture, the selected split points, and the number of layers used in this work serve as a specific example corresponding to the evaluation presented in Section V. In practice, the neural network designer must analyze the target model to identify appropriate split points and determine the number of additional layers to execute after each truncation. These decisions depend on the network’s structure and the desired balance between accuracy and computational efficiency under dynamic conditions.
V Performance Evaluation
Our simulation environment is designed to emulate the dynamic conditions of a large-scale UAV swarm executing split computing tasks. In our environment, UAVs are deployed over a 20×20 km area and follow circular trajectories with a movement radius of 1000 m, traveling at speeds of up to 75 m/s. Task arrivals are modeled as a Markov process with an average inter-arrival time of 60 ms, which accurately captures the random and bursty nature of workload generation. In addition, the simulator captures detailed communication and computation aspects—including realistic channel characteristics, transmission delays, and the heterogeneous computational capabilities of each UAV ensuring that both the physical movement and the fluctuating workload distribution among nodes are properly considered. Table I shows the rest of the simulation parameters.
Within this dynamic simulation framework, we evaluate five distinct task offloading strategies. In the first scenario, tasks are offloaded in a purely random manner to a neighboring UAV, providing a basic baseline for performance comparison. The second strategy follows a similar random selection process, but enforces an acyclic constraint to prevent redundant offloading cycles. The third approach adopts a greedy policy, where each UAV offloads its task to the least busy neighbor based on instantaneous load measurements, thus aiming to minimize processing delays. The fourth strategy, referred to as LocalOnly, eliminates offloading entirely so that tasks are processed solely on the originating UAV, which highlights the trade-off between local processing overhead and the benefits of collaborative offloading. Finally, our proposed Distributed method leverages a diffusive metric—in particular, the aggregated gigaflops measure—to dynamically route tasks based on both the local UAV’s capacity and the capabilities of its neighbors. This comprehensive set of cases allows us to rigorously assess how different offloading mechanisms perform under the challenging conditions of a mobile, Markov-driven UAV swarm.
In our experiments, we consider the neural network model described in Section III pre-trained offline and deployed across the UAV swarm. The full model achieves an accuracy of 0.95, while the early-exit configurations at the 30th and 15th layers provide accuracies of 0.9 and 0.6, respectively. These values quantify the trade-off between computational savings and inference quality when applying early exits under high-load conditions. Lastly, the threshold $\gamma$ for distributed allocation is set to 0.02 based on our empirical data gathered from the empirical simulations. We selected this value to effectively balance the trade-off between frequent task transfers and achieving equitable workload distribution.
TABLE I: Simulation Parameters
@@ V Performance Evaluation
V-A Basic Simulations
@@
V-B Effect of Input Rate
@@
Figure 4: Performance metrics under varying task arrival rates for 30 workers. Panel (a) shows average latency, (b) displays remaining GFLOPs, and (c) presents the overall figure of merit.
V-C Effect of Network Distribution

### v2

@@ 1 Introduction
The escalating prominence of Unmanned Aerial Vehicle (UAV) swarms has opened up new horizons in applications spanning surveillance, disaster response, smart agriculture, and beyond. By coordinating large numbers of autonomous aerial units, system-level benefits such as enhanced coverage, redundancy, and adaptability can be realized. However, these advantages come hand-in-hand with steep technical challenges, particularly when it comes to designing and deploying machine learning (ML) services in vast swarms. Traditional methods that rely on a single, centralized server or controller become increasingly untenable as the number of nodes and geographic scale increase [distributed_survey_1]. A centralized ground station, faces prohibitive communication overheads, latency bottlenecks, and reliability issues. Moreover, real-world UAV swarms are dynamic, meaning that nodes may join, leave, or rearrange themselves in the network. Such dynamic conditions only make the limitations of traditional centralized paradigms worse [islam2021survey].
@@
Recent efforts to alleviate centralized burdens typically advocate a form of distributed operation by distributing inference or training tasks across multiple nodes [distributed_survey_2]. However, they still often rely on global awareness or explicit synchronization. At smaller scales or in constrained network topologies, these partial-distribution schemes might suffice. Yet, as UAV swarms expand to hundreds or even thousands of nodes, it becomes infeasible to gather and coordinate global network information [Tahir2019SwarmsOU]. Additionally, heterogeneous node capabilities such as differences in on-board processing, sensing, and power resources demand an approach that can dynamically adjust to varying node states in near real-time. A node endowed with high computation power a time may deplete its battery and become less capable a moment later, or an agile UAV that was ideally placed to handle tasks at one moment could relocate out of range in the next. These variations highlight the necessity for a self-organizing approach that pushes intelligence and decision-making to the edges.
@@
Split learning and computing is particularly well-suited to address these challenges in large-scale UAV swarms [distributed_survey_2]. Rather than transmitting raw data or entire models to a single aggregator, split learning and computing partitions the model into layers distributed among various nodes, thereby enabling partial inference at local levels. By transferring only the necessary, intermediate representations, nodes can reduce communication overhead while maintaining data privacy. However, effectively implementing split computing in UAV networks requires strategic decision-making about when and where to forward intermediate layer outputs [distributed_survey_1, distributed_survey_2]. This is where centralized schemes fail to perform well for large networks. They attempt to collect global knowledge of the network’s topology, node resource levels, and data distributions which is impractical in large, dynamic swarms. As a result, model splits and partial inferences often end up bottlenecked by nodes that unknowingly become overloaded or unreachable.
@@
To further enhance the system’s adaptability under fluctuating workloads, we incorporate an early-exit mechanism that dynamically reduces the depth of inference based on real-time load conditions. This strategy allows nodes to truncate the model execution at predefined intermediate layers, producing acceptable outputs while avoiding the full computational cost of processing the entire model. By adjusting the exit point according to the current resource availability and task queue length, the system minimizes latency and prevents resource saturation. This approach aligns with the principles of early exiting  [matsubara2023split] where intermediate outputs offer a controllable trade-off between computational effort and inference accuracy, making it particularly suitable for dynamic and resource-constrained UAV environments.
@@
The remainder of this paper is organized as follows. Section 2 reviews related work on distributed inference and task offloading in UAV networks. Section 3 formalizes the system model and problem definition, introducing the proposed diffusive metric and task allocation strategy. Section 5 presents the simulation setup and provides a detailed performance evaluation under various operating conditions. Finally, Section 6 concludes the paper and outlines directions for future work.
@@ 2 Related Work
Recent research has explored distributed DNN inference and task offloading to address the challenges posed by limited resources and dynamic network conditions. In  [modnn], MoDNN partitions DNN computations among mobile devices to balance workload and reduce communication overhead; its model partitioning strategy minimizes transmission delays while exploiting parallel processing, making it highly applicable to resource-constrained environments. Similarly, DISNET  [disnet] introduces a micro-split framework that partitions deep neural network models both vertically and horizontally, enabling cooperative inference among heterogeneous IoT devices while reducing latency and energy consumption through adaptive workload allocation.
@@
On the split computing front,  [Cao2024Multitier] propose a multi tier framework that employs deep reinforcement learning to optimize model splitting and resource allocation in UE–Edge–Cloud systems, decomposing complex optimization problems into sequential subproblems to balance computational and communication loads. In a similar vein, CoEdge  [zeng2020] demonstrates that dynamic workload partitioning, formulated as a constrained optimization problem, can yield substantial latency speedups, with a fast approximation algorithm that adapts in real time to varying device capabilities and network conditions. Similarly for Split Learning,  [lin2024epsl] develop an efficient parallel split learning (EPSL) framework for resource-constrained wireless edge networks. Their approach jointly optimizes subchannel allocation, transmit power, and model split point selection to minimize latency and communication overhead.  [kim2024edgemsl] propose Edge-MSL, a split learning framework for mobile edge networks that leverages a contextual multi-armed bandit formulation to dynamically select the optimal split point and assign edge servers. By adapting the split location and offloading decisions based on real-time network conditions, their approach significantly reduces training latency and enhances convergence. On the other hand,  [hybridsplit] combines split learning with federated learning in UAV networks to address the trade-off between communication overhead and computational constraints, resulting in an architecture that achieves higher inference accuracy than standalone federated learning while reducing transmission costs compared to conventional split learning. Moreover,  [wu2023] propose a parallel split learning framework over wireless networks that integrates dynamic resource management with parallel task offloading. Their method exploits concurrent processing across heterogeneous edge devices to substantially reduce training latency while maintaining balanced workload distribution.  [yeom2024uav] investigate a UAV-assisted split computing system where a UAV acts as a mobile edge server to support collaborative inference among IoT devices. By partitioning a DNN between on-device processing and UAV-based computation, their design reduces transmission delays and energy consumption.
@@
Task offloading approaches can also be useful for split learning approaches. For instance,  [fei] develops a UAV-based environmental monitoring system that employs on-board processing to minimize transmission delays and maintain low data freshness latency, thereby enhancing overall system capacity in remote areas.  [cloud_mec] proposes a cloud–MEC collaborative offloading scheme that integrates differentiated offloading decisions with service orchestration via an ODaS mechanism, significantly reducing both delay and energy consumption by balancing computational and communication loads.  [zhou2023dag] addresses the challenge of offloading dependent tasks modeled as directed acyclic graphs using a branch soft actor-critic algorithm, which optimizes offloading, computation, and cooperation decisions under dynamic conditions to reduce latency and improve energy efficiency. More recently,  [zheng2024multi] presents a multi-agent collaborative framework that jointly optimizes UAV trajectory planning and DAG task offloading via deep reinforcement learning, with their TD3-TT algorithm demonstrating further latency reductions and enhanced real-time performance. The FOCUS framework  [focus] leverages fog computing in UAV networks by jointly optimizing routing and computation. Its central SDN-based architecture and analytical approach achieve significant latency improvements by effectively integrating on-board processing with data forwarding. Lastly,  [agarwal2024federated] investigate task offloading in a UAV-assisted MEC network by leveraging a federated learning framework. In their work, a UAV equipped with MEC capabilities serves as a mobile server to assist resource-constrained UE in offloading computation-intensive tasks.
@@
3 Problem Formulation and System Model
@@
3.1 Task Computation Model
@@ 3.2 Communications
3.3 Energy Consumption Model
@@
4 Distributed Task Allocation and Early-Exit Mechanism
@@
Nodes with powerful neighbors or central positions are more advantageous for task allocation. To identify such nodes, we aim to utilize graph centrality measures. Metrics like eigenvector centrality, which propagate importance through neighbors, appear suitable for capturing this effect. However, computing eigenvector centrality requires global normalization at every iteration, making it impractical for a fully distributed and rapidly changing UAV network.
@@
To address this issue, we propose using aggregated computation capability instead of eigenvector centrality. This metric preserves the neighbor-based recursive structure while removing the need for global normalization. Eigenvector centrality updates each node using the contributions of its neighbors, and our formulation mirrors this idea: each node’s effective computation capability is updated based on its own resources and the aggregated capability of its neighbors.
We compute aggregated computation capability iteratively. The intuition is that when a node splits an incoming load evenly among itself and its neighbors, the total completion time is dictated by the slowest participant. This leads to the following neighbor-recursive update for the aggregated computation capability $\phi_{i,t+1}$ of node $v_{i}$ at time $t+1$:
@@
where $M_{i}(t)$ is the neighbor set of node $i$, $F_{i}$ is computation capability of node $i$, $d^{tx}_{i,k}(t)$ is the transmission delay between nodes $i$ and $k$, and $\phi_{i,t}$ is the aggregated computation capability of node $i$ and time $t$. The aggregated computation capability assumes that a given amount of work is divided evenly among the node and its neighbors ($|M_{i}(t)|+1$). The node processes its own share and forwards the remaining portions to its neighbors. Since all participants work concurrently, the overall completion time is determined by the slowest node, whose aggregated capability is captured by the maximum term. Because this update requires only the neighbors of the nodes, it naturally adapts to dynamic network conditions without requiring any global view.
4.1 Task Transfer Decision
Our framework requires each UAV to make autonomous decisions about whether to process a task locally or transfer it to a neighboring node. These decisions are based on comparative utilization levels between nodes, mediated by a threshold parameter $\gamma$ that controls the sensitivity of task offloading. The utilization of node $i$ at time $t$ is calculated as the ratio of its current computational load to its aggregated computation capability:
@@ 4.1 Task Transfer Decision
4.2 Early-Exit Mechanism
@@
This adaptive mechanism ensures that nodes under high load degrade computation gracefully, trading off accuracy to maintain responsiveness and system stability. It is important to note that the neural network architecture, the selected split points, and the number of layers used in this work serve as a specific example corresponding to the evaluation presented in Section 5. In practice, the neural network designer must analyze the target model to identify appropriate split points and determine the number of additional layers to execute after each truncation. These decisions depend on the network’s structure and the desired balance between accuracy and computational efficiency under dynamic conditions.
Table 1: Simulation Parameters
@@
5 Performance Evaluation
Our simulation environment is designed to emulate the dynamic conditions of a large-scale UAV swarm executing split computing tasks. In our environment, UAVs are deployed over a 20×20 km area and follow circular trajectories with a movement radius of 1000 m, traveling at speeds of up to 75 m/s. Task arrivals are modeled as a Markov process with an average inter-arrival time of 60 ms, which accurately captures the random and bursty nature of workload generation. In addition, the simulator captures detailed communication and computation aspects,including realistic channel characteristics, transmission delays, and the heterogeneous computational capabilities of each UAV ensuring that both the physical movement and the fluctuating workload distribution among nodes are properly considered. Table 1 shows the rest of the simulation parameters.
Within this dynamic simulation framework, we evaluate five distinct task offloading strategies. In the first scenario, tasks are offloaded in a purely random manner to a neighboring UAV, providing a basic baseline for performance comparison. The second strategy follows a similar random selection process, but enforces an acyclic constraint to prevent redundant offloading cycles. The third approach adopts a greedy policy, where each UAV offloads its task to the least busy neighbor based on instantaneous load measurements, thus aiming to minimize processing delays. The fourth strategy, referred to as LocalOnly, eliminates offloading entirely so that tasks are processed solely on the originating UAV, which highlights the trade-off between local processing overhead and the benefits of collaborative offloading. Finally, our proposed Distributed method leverages a diffusive metric, to dynamically route tasks based on both the local UAV’s capacity and the capabilities of its neighbors. This comprehensive set of cases allows us to rigorously assess how different offloading mechanisms perform under the challenging conditions of a mobile, Markov-driven UAV swarm.
In our experiments, we consider the neural network model described in Section 3 pre-trained offline and deployed across the UAV swarm. The full model achieves an accuracy of 0.95, while the early-exit configurations at the 30th and 15th layers provide accuracies of 0.9 and 0.6, respectively. These values quantify the trade-off between computational savings and inference quality when applying early exits under high-load conditions. Lastly, the threshold $\gamma$ for distributed allocation is set to 0.02 based on our empirical data gathered from the empirical simulations. We selected this value to effectively balance the trade-off between frequent task transfers and achieving equitable workload distribution.
5.1 Basic Simulations
@@
Figure 4: Performance metrics under varying task arrival rates for 30 workers. Figure (a) shows average latency, (b) displays remaining GFLOPs, and (c) presents the overall figure of merit.
Figure 5: Performance metrics under varying mission area sizes. Figure (a) shows average latency, (b) depicts remaining GFLOPs, and (c) presents the overall figure of merit.
5.2 Effect of Input Rate
@@
5.3 Effect of Network Distribution